OpenStudy (anonymous):

The Community Recreation Center is developing plans for a new sports facility. Community members can submit suggestions for the new facility, along with basic scale drawings of their ideas. Elena wants to include a new 15- by 30-meter basketball court in the athletic center. She is submitting a scale drawing on an 8.5- by 11-inch sheet of paper. Which scale should Elena use to create as large a drawing as possible on the paper? A. 1/2 in. = 1 m B. 5/16 in. = 1 m C. 1/4 in. = 1 m D. 3/8 in. = 1 m

OpenStudy (uglyturtle):

The Community Center is 15 by 30 meters. The ratio is 15:30 which equals 1:2 or 1/2. It seems that the most accurate representation would be 1/2 in. = 1 m.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (uglyturtle):

I will try to.

OpenStudy (anonymous):

The Community Recreation Center is developing plans for a new sports facility. Community members can submit suggestions for the new facility, along with basic scale drawings of their ideas. Lupe wants to include a new 65- by 100-meter soccer field in the athletic center. She is submitting a scale drawing on an 8.5- by 11-inch sheet of paper. Which scale should Lupe use to create as large a drawing as possible on the paper?

OpenStudy (uglyturtle):

65/100 = 13/20

OpenStudy (anonymous):

A. 1/4 in. = 1 m B. 5/8 in. = 1 m C. 1/8 in. = 1 m D. 1/16 in. = 1 m

OpenStudy (uglyturtle):

What are the options for the answers?

OpenStudy (uglyturtle):

Okay.

OpenStudy (uglyturtle):

Well this seems to be a comparison of fractions.

OpenStudy (uglyturtle):

And 100/8 = 12.5 > 11

OpenStudy (uglyturtle):

So it has to be 1/16.

OpenStudy (uglyturtle):

Goodnight.

OpenStudy (anonymous):

night